Circular No. 6489

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

V2116 OPHIUCHI F. Jablonski, Laboratorio Nacional de Astrofisica (LNA), Itajubá; M.

Pereira and J. Braga, Instituto Nacional de Pesquisas Espaciais, S.J. Campos, Brazil; and C. Gneiding, LNA, report on optical observations of V2116 Oph, the counterpart of the accreting x-ray pulsar GX 1+4, carried out during Apr.–Aug. with the LNA 1.6-m and 0.6-m telescopes (+ CCD fast photometer): “On April 26, from 3300 images (integration time 5 s each) at the 1.6-m telescope, sinusoidal optical pulsations close to the ∼ 2-min x-ray period were clearly detected for the first time in V2116 Oph. The barycentric pulse period was 124.171 ± 0.004 s and the semiamplitude of the pulsations was 1.3 ± 0.1 percent. The passband for the observations was defined by a CuSO_{4} filter. The power spectrum of the light curve shows a monochromatic signal with peak power ∼ 220 times the local continuum. None of six comparison stars (measured simultaneously) shows any feature with significant spectral power. Follow-up observations in different conditions (with other filters and detectors) confirmed the previous detections, showing relative semiamplitudes of the pulsations of up to ∼ 4 percent. The pulse period is consistent with BATSE and RXTE results (IAUC 6478, 6153) projected to the epoch of the observations. There is a definite correlation between detectability of pulsations and the optical brightness of V2116 Oph. No pulsations were detected when the system was fainter than R ∼ 17. Flickering similar to that observed in cataclysmic variables is seen when V2116 Oph is bright. Examples of light curves obtained with LNA’s fast photometer can

be seen on the Web at http://www.lna.br/instrum/novida e.html.”

COMET C/1996 R1 (HERGENROTHER-SPAHR) Ephemeris continuation to IAUC 6479 (elements MPC 27882):

1996TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Oct. 14 20^{h}58.^{m}50 − 2^{o}50.^{′}3 1.352 1.985 114.^{o}5 27.^{o}2 12.6
24 20 42.24 − 9 31.3 1.608 2.024  99.4 29.0 13.1
Nov. 3 20 33.37 −14 06.2 1.882 2.069  86.2 28.6 13.5
13 20 29.41 −17 19.6 2.159 2.120  74.5 26.7 13.9
23 20 28.79 −19 40.2 2.426 2.176  63.7 24.0 14.3
Dec. 3 20 30.46 −21 26.1 2.678 2.237  53.5 20.8 14.6
13 20 33.76 −22 48.7 2.908 2.302  43.9 17.2 14.9
23 20 38.19 −23 55.6 3.113 2.371  34.6 13.6 15.2
1996 October 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green